OpenStudy (anonymous):

how do you find 25% of 96?

OpenStudy (anonymous):

25% = 25/100 \[\text{25% of M} = {25 \over 100} \times M \]

OpenStudy (anonymous):

So 25% of 96 = ?

OpenStudy (anonymous):

ummm explain please

OpenStudy (anonymous):

24

OpenStudy (anonymous):

I just did! You explain what you don't understand. Yes, 24 is correct.

OpenStudy (anonymous):

A percent is just a number divided by 100. \[P\ \% = {P \over 100}\]

OpenStudy (anonymous):

And typically 'of' means multiply.